# On the stability of global forecasting models

### v1

@@ 4 Results and discussion
Figure 6: MQC results for each ensemble and retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$ for the M5 dataset and $r=1$ for the VN1 dataset.
Figures 7 and 8 illustrate the fundamental trade-off between forecast accuracy and stability, comparing RMSSE with sMAPC and MQL with MQC, respectively. These plots reveal that high accuracy does not guarantee high stability. In particular, several deep learning models, achieve low RMSSE or MQL scores but exhibit high instability, especially in VN1, highlighting their sensitivity to changes in the training data and their volatility across forecast origins. Conversely, models like Random Forest, XGBoost, and LGBM strike a more favorable balance, maintaining relatively low RMSSE and sMAPC values. Nevertheless, it is the ensemble models, and especially those with larger size like Ens4A and Ens5A, that most consistently approach the optimality of stability and accuracy. In both the point and probabilistic domains, even very simple ensembles manage to control variance and mitigate instability without sacrificing much, if any, predictive performance. The overall picture that emerges from these figures reinforces the idea ensembling offers a practical and effective strategy to counteract forecast instability, particularly when constructed from a diverse pool of models.
Figure 7: The accuracy-stability trade-off. RMSSE vs sMAPC for the M5 and VN1 datasets.
Figure 8: The accuracy-stability trade-off. MQL vs MQC for the M5 and VN1 datasets.
This set of results has clear practical implications: in operational forecasting systems, especially in domains such as retail and supply chain management, the use of ensembles combined with lower retraining frequencies may offer a robust forecasting process, balancing performance with reliability.
@@ 5 Conclusions
This study provides a comprehensive empirical assessment of forecast stability in global models, focusing on the effects of retraining frequency and ensembling across both point and probabilistic forecasting contexts. Our findings demonstrate that global forecasting models are able to produce very stable predictions and that they can achieve even more stability under reduced retraining frequencies, challenging the prevailing wisdom that frequent model updates are necessary to maintain forecast reliability.
In terms of point forecasting, the results consistently show that retraining less frequently does not harm, and often improves, forecast stability. This trend is particularly evident in the M5 dataset, where many models exhibit monotonic or flat stability profiles as the retraining interval increases. Similar patterns, although more nuanced, are observed in the VN1 dataset. Probabilistic forecast stability, as measured by the newly defined Multi-Quantile Change (MQC), shows a slightly different behavior. While in the M5 dataset forecast stability remains largely unaffected by retraining frequency, in VN1 we observe a convex relationship, indicating that retraining up to every 10 weeks can preserve stability, but beyond that point the quality of probabilistic forecasts begins to deteriorate. These results are statistically validated via Friedman-Nemenyi tests and indicate that continuous retraining is often not the optimal strategy.
Ensembling emerges as a powerful strategy to enhance forecast stability. While combining only the most accurate models offers moderate gains, the most substantial improvements in stability, both point and probabilistic, are achieved by ensembles that incorporate a diverse mix of forecasting algorithms. This effect is especially pronounced in the VN1 dataset, where larger and more heterogeneous ensembles significantly reduce forecast volatility across retraining scenarios. Notably, ensemble models not only improve stability metrics but also smooth the variation in stability across retraining frequencies, further reinforcing their role as a stabilizing force in global forecasting systems. These results suggest that model diversity should be prioritized over raw model performance when building ensembles.
@@
From a practical standpoint, these findings have several implications. First, businesses that rely on large-scale forecasting systems, such as those in retail, inventory management, and supply chain planning, can adopt lower retraining frequencies not only to conserve computational resources but also to increase operational stability, without compromising on forecast quality. Second, incorporating diverse ensembles into global forecasting frameworks provides an effective hedge against model instability, thereby increasing the reliability of forecasts used in decision-making processes. Finally, these results support a shift in emphasis from pure accuracy to stability-aware model selection and evaluation, in those contexts that are more sensitive to forecast reviews.
Nevertheless, some limitations of the present study should be acknowledged. While the analysis covers a broad range of global models and two state-of-the-art retail forecasting datasets, the generalizability of the findings to other domains or data types (e.g., macroeconomic indicators, financial series, or healthcare time series) remains an open question. Additionally, the study focuses solely on full retraining schemes and does not explore incremental or online learning strategies, which may offer a different trade-off between accuracy and stability. Future research could extend this work by incorporating such adaptive methods, assessing alternative ensembling techniques (e.g., weighted or stacking ensembles), and introducing business-oriented stability metrics that link forecast instability directly to operational cost or risk. Further theoretical development of stability measures, particularly for probabilistic forecasts, may also contribute to a deeper understanding of how models behave across time and under varying data regimes.
In conclusion, this work provides strong empirical evidence that frequent retraining is not necessary for maintaining stability in global forecasting models, and that ensembling, particularly when built on diverse model pools, is a highly effective strategy for achieving both stable and accurate forecasts. These findings contribute to a growing body of research promoting sustainable and robust forecasting practices in modern data-intensive environments.
@@
In this section, we provide tables and figures related to the empirical results of the M5 and VN1 datasets. Table 4 shows the composition of ensembles across the different datasets.
@@
The Tables 5, 6, 8, and 7 show the forecast instability of the different models and ensembles along the examined retrain scenarios for the M5 daily dataset.
@@
LR 0.118 0.118 0.118 0.119 0.119 0.119 0.119 0.119 0.119 0.119
XGBoost 0.115 0.115 0.115 0.116 0.115 0.115 0.115 0.116 0.115 0.114
LGBM 0.123 0.122 0.123 0.124 0.123 0.123 0.123 0.124 0.123 0.122
CatBoost 0.127 0.144 0.123 0.144 0.144 0.144 0.144 0.142 0.141 0.139
MLP 0.109 0.109 0.110 0.110 0.111 0.111 0.112 0.111 0.110 0.110
TCN 0.114 0.114 0.114 0.113 0.113 0.112 0.113 0.112 0.110 0.108
NBEATSx 0.110 0.109 0.110 0.115 0.110 0.112 0.111 0.111 0.129 0.169
NHITS 0.112 0.111 0.122 0.112 0.114 0.114 0.115 0.135 0.114 0.115
Table 7: M5 MQC values for each method and retrain scenario combination.
@@
Ens2A 0.117 0.117 0.117 0.118 0.118 0.118 0.118 0.119 0.118 0.117
Ens3A 0.117 0.117 0.116 0.118 0.118 0.118 0.117 0.118 0.117 0.117
Ens4A 0.113 0.113 0.113 0.114 0.114 0.113 0.113 0.114 0.113 0.113
Ens5A 0.111 0.111 0.111 0.112 0.111 0.111 0.111 0.111 0.110 0.110
Table 8: M5 MQC values for each ensemble and retrain scenario combination.
The Tables 9, 10, 12, and 11 show the forecast instability of the different models and ensembles along the examined retrain scenarios for the VN1 weekly dataset.
@@
LR 1.064 1.067 1.063 1.061 1.070 1.079 1.103 1.097 1.127 1.163
RF 1.322 1.329 1.328 1.341 1.353 1.366 1.403 1.392 1.439 1.504
XGBoost 1.101 1.072 1.080 1.079 1.083 1.097 1.084 1.135 1.175 1.254
LGBM 1.282 1.179 1.235 1.161 1.165 1.173 1.199 1.277 1.266 1.379
CatBoost 1.435 1.281 1.244 1.159 1.150 1.166 1.151 1.225 1.207 1.246
MLP 0.844 0.876 0.887 0.883 0.889 0.891 0.876 0.933 0.903 0.969
LSTM 1.020 1.015 1.008 1.004 1.005 0.993 0.986 0.987 0.972 0.930
TCN 1.020 1.015 1.008 1.004 1.005 0.993 0.986 0.987 0.972 0.930
NBEATSx 1.532 1.519 1.467 1.253 1.167 1.210 1.196 1.069 1.084 1.147
NHITS 1.592 1.505 1.412 1.254 1.212 0.998 1.174 1.028 1.082 1.145
Table 11: VN1 MQC values for each method and retrain scenario combination.
@@
Ens2A 1.079 1.114 1.062 0.974 0.949 0.982 0.962 0.921 0.944 1.022
Ens3A 1.200 1.192 1.148 1.024 0.990 0.973 0.995 1.005 0.970 1.034
Ens4A 1.180 1.166 1.122 1.036 1.021 1.001 1.028 0.988 1.017 1.080
Ens5A 1.153 1.137 1.107 1.033 1.022 1.000 1.031 1.006 1.029 1.091
Table 12: VN1 MQC values for each ensemble and retrain scenario combination.
Figure 9: M5 Friedman-Nemenyi test results based on MQC.
Figure 10: VN1 Friedman-Nemenyi test results based on MQC.
@@
Figure 12: M5 Friedman-Nemenyi test results on ensembles based on MQC.
@@
Figure 14: VN1 Friedman-Nemenyi test results on ensembles based on MQC.

### v2

@@ 4 Results and discussion
Figure 6: SMQC results for each ensemble and retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$ for the M5 dataset and $r=1$ for the VN1 dataset.
Figures 7 and 8 illustrate the fundamental trade-off between forecast accuracy and stability, comparing RMSSE with sMAPC and SMQL with SMQC, respectively. These plots reveal that high accuracy does not always guarantee high stability. In particular, several models achieve low RMSSE or SMQL scores but exhibit high instability, especially in VN1, highlighting their sensitivity to changes in the training data and their volatility across forecast origins. Nevertheless, it is the ensemble models that most consistently approach the optimality of stability and accuracy. In both the point and probabilistic domains, even very simple ensembles manage to control variance and mitigate instability without sacrificing much, if any, predictive performance. The overall picture that emerges from these figures reinforces the idea ensembling offers a practical and effective strategy to counteract the forecast instability.
Figure 7: The accuracy-stability trade-off. RMSSE vs sMAPC for the baseline scenario, $r=7$ for the M5 dataset, and $r=1$ for the VN1 dataset.
Figure 8: The accuracy-stability trade-off. SMQL vs SMQC for the baseline scenario, $r=7$ for the M5 dataset, and $r=1$ for the VN1 dataset.
This set of results has clear practical implications: in operational forecasting systems, especially in domains such as retail and supply chain management, the use of ensembles combined with lower retraining scenarios may offer a robust forecasting process, balancing performance with stability.
@@ 5 Conclusions
This study provides a comprehensive empirical assessment of the forecast stability of global models, focusing on the effects of retraining across both point and probabilistic forecasting, in the context of retail demand. Our findings demonstrate that global forecasting models are able to produce very stable predictions, and that they can achieve even more stability under reduced retraining strategies, challenging the prevailing wisdom that frequent model updates are necessary to maintain forecast stability.
In terms of point forecasting, the results consistently show that retraining less frequently does not harm and often improves forecast stability. This trend is evident in both the M5 and the VN1 datasets, where many models exhibit monotonic or flat stability profiles as the retraining interval increases. Probabilistic forecast stability, as measured by the newly defined Scaled Multi-Quantile Change (SMQC), shows a very similar behavior, indicating that the stability of probabilistic forecasts does not deteriorate as the models are retrained less frequently. These results are statistically validated via Friedman-Nemenyi tests and indicate that continuous retraining is often not the optimal strategy.
Ensembling can be considered a good strategy to enhance forecast stability. Notably, ensemble models not only improve stability metrics but also smooth the variation in stability across retraining scenarios, further reinforcing their role as a stabilizing force in global forecasting systems. However, from our analysis, it is not clear whether the size and the diversity of the ensemble may have an impact on stability, and more research is needed in this direction.
@@
From a practical standpoint, these findings have several implications. First, businesses that rely on large-scale forecasting systems, such as those in retail, inventory management, and supply chain planning can adopt lower retraining strategies not only to conserve computational resources, but also to increase operational stability, without compromising on forecast accuracy. Second, incorporating ensembles into global forecasting frameworks provides an effective hedge against model instability, thereby increasing the reliability of forecasts used in decision-making processes. Finally, these results support a shift in emphasis from pure accuracy to stability-aware model selection and evaluation, in those contexts that are more sensitive to forecast reviews.
Nevertheless, some limitations of the present study should be acknowledged. While the analysis covers a broad range of global models and two state-of-the-art retail forecasting datasets, the generalizability of the findings to other domains or data types (e.g., macroeconomic indicators, financial series, or healthcare time series) remains an open question. Moreover, this work deliberately focuses on the stability of global models, leaving the analysis of stability in local models as an open avenue. A direct comparison between global and local approaches could therefore represent a valuable direction for future research. Additionally, the study focuses solely on full retraining schemes and does not explore incremental or online learning strategies, which may offer a different trade-off between accuracy and stability. Future research could extend this work by incorporating such adaptive methods, assessing alternative ensembling techniques (e.g., weighted or stacking ensembles), and introducing business-oriented stability metrics that link forecast instability directly to operational cost or risk. Furthermore, this study employed a single framework for generating probabilistic forecasts, namely Conformal Inference. While conformalized residuals provide a valid means for quantifying uncertainty, the resulting quantiles may primarily capture the stability of the conformal adjustment rather than that of the underlying predictive distributions. Future research could address this limitation by studying probabilistic stability in native probabilistic models, thereby enhancing the robustness and generalizability of the findings. Finally, further theoretical development of stability measures, particularly for probabilistic forecasts, may also contribute to a deeper understanding of how models behave across time and under varying data regimes.
In conclusion, this work provides strong empirical evidence that frequent retraining is not necessary for maintaining stability in global forecasting models, and that ensembling is an effective strategy for achieving both stable and accurate forecasts. These findings contribute to a growing body of research promoting sustainable and robust forecasting practices in modern data-intensive environments.
@@
In this section, we provide tables and figures related to the empirical results of the M5 and VN1 datasets.
Table 3 shows the composition of ensembles across the different datasets.
The Tables 4, 5, 7, and 6 show the forecast instability of the different models and ensembles along the examined retrain scenarios for the M5 daily dataset.
The Tables 8, 9, 11, and 10 show the forecast instability of the different models and ensembles along the examined retrain scenarios for the VN1 weekly dataset.
@@
LR 0.110 0.110 0.110 0.111 0.111 0.111 0.112 0.112 0.112 0.113
XGBoost 0.100 0.099 0.099 0.100 0.100 0.100 0.099 0.100 0.099 0.098
LGBM 0.118 0.117 0.117 0.119 0.118 0.118 0.118 0.119 0.118 0.118
CatBoost 0.156 0.157 0.154 0.157 0.157 0.157 0.157 0.156 0.156 0.155
MLP 0.083 0.083 0.083 0.083 0.084 0.084 0.085 0.085 0.083 0.085
TCN 0.083 0.083 0.083 0.082 0.082 0.082 0.082 0.082 0.080 0.080
NBEATSx 0.084 0.084 0.084 0.088 0.085 0.085 0.085 0.086 0.088 0.105
NHITS 0.087 0.085 0.101 0.086 0.087 0.088 0.088 0.091 0.088 0.089
Table 6: M5 SMQC values for each method and retrain scenario combination.
@@
Ens2A 0.108 0.108 0.108 0.108 0.108 0.108 0.108 0.108 0.108 0.107
Ens3A 0.107 0.107 0.107 0.108 0.107 0.107 0.107 0.108 0.107 0.107
Ens4A 0.099 0.099 0.099 0.101 0.100 0.100 0.099 0.100 0.099 0.099
Ens5A 0.095 0.095 0.095 0.096 0.096 0.095 0.095 0.096 0.095 0.095
Table 7: M5 SMQC values for each ensemble and retrain scenario combination.
@@
LR 1.560 1.552 1.546 1.543 1.562 1.525 1.547 1.643 1.682 1.900
RF 0.729 0.732 0.735 0.737 0.735 0.738 0.743 0.734 0.731 0.734
XGBoost 0.629 0.598 0.600 0.584 0.585 0.575 0.556 0.613 0.599 0.589
LGBM 1.484 1.433 1.457 1.471 1.502 1.349 1.459 1.394 1.593 1.445
CatBoost 2.276 1.861 1.769 1.582 1.547 1.579 1.333 1.581 1.302 1.243
MLP 0.219 0.281 0.278 0.278 0.282 0.275 0.266 0.285 0.267 0.283
LSTM 0.242 0.238 0.237 0.235 0.233 0.229 0.227 0.224 0.214 0.187
TCN 0.241 0.238 0.237 0.235 0.233 0.229 0.227 0.224 0.214 0.187
NBEATSx 0.765 0.763 0.708 0.567 0.477 0.442 0.520 0.348 0.362 0.369
NHITS 0.755 0.779 0.660 0.555 0.506 0.333 0.507 0.348 0.343 0.365
Table 10: VN1 SMQC values for each method and retrain scenario combination.
@@
Ens2A 0.441 0.476 0.452 0.391 0.354 0.332 0.364 0.300 0.299 0.312
Ens3A 0.539 0.562 0.507 0.430 0.389 0.324 0.398 0.312 0.309 0.322
Ens4A 0.545 0.557 0.522 0.465 0.433 0.387 0.443 0.376 0.376 0.389
Ens5A 0.540 0.545 0.519 0.472 0.448 0.410 0.453 0.407 0.405 0.416
Table 11: VN1 SMQC values for each ensemble and retrain scenario combination.
Figure 9: M5 Friedman-Nemenyi test results based on SMQC.
Figure 10: VN1 Friedman-Nemenyi test results based on SMQC.
@@
Figure 12: M5 Friedman-Nemenyi test results on ensembles based on SMQC.
@@
Figure 14: VN1 Friedman-Nemenyi test results on ensembles based on SMQC.